Composition of the nuclei of two atomic species X and Y are given as under
                      X           Y
Protons  =      6            6
Neutrons =      6            8 
         Give the mass number of X and Y. What is the relation between the two species?

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Mass number of X = Protons + Neutrons
= 6 + 6
= 12
Mass number of Y = Protons + Neutrons
= 6 + 8 = 14
As the atomic number is same i.e., = 6
[atomic number = number of protons].
Both X and Y are isotopes of same element.

Thanks